eEgg is considering the purchase of a new distributed network computer system to help handle its warehouse inventories. The system costs $40,000 to purchase and install and $29,000 to operate each year. The system is estimated to be useful for 4 years. Management expects the new system to reduce the cost of managing inventories by $50,000 per year. The firm’s cost of capital (discount rate) is 9%.

Required:

1. What is the net present value (NPV) of the proposed investment under each of the following independent situations? (Use the appropriate present value factors from Appendix C, TABLE 1 and Appendix C, TABLE 2.)

1a. The firm is not yet profitable and therefore pays no income taxes.

1b. The firm is in the 27% income tax bracket and uses straight-line (SLN) depreciation with no salvage value. Assume MACRS rules do not apply.

1c. The firm is in the 27% income tax bracket and uses double-declining-balance (DDB) depreciation with no salvage value. Given a four-year life, the DDB depreciation rate is 50% (i.e., 2 × 25%). In year four, record depreciation expense as the net book value (NBV) of the asset at the start of the year.

2. What is the internal rate of return (IRR) of the proposed investment for situations in requirement 1, parts (a) through (c)? Use the IRR builit-in function in Excel to compute the IRR.
Answer #1
Initial Investment $40,000
1a
Year 0 1 2 3 4
Initial Investment -$40,000
Operating cost -$29,000 -$29,000 -$29,000 -$29,000
Saving in cost $50,000 $50,000 $50,000 $50,000
Total CashFlow -$40,000 $21,000 $21,000 $21,000 $21,000
PV Factor @ 9%= (1/(1.09^n)) 1 0.9174 0.8417 0.7722 0.7084
NPV -$40,000 $19,266 $17,675 $16,216 $14,877 $28,034
IRR (using IRR Formula) 38.04%
1b
SLM Dep. = (Cost-salvage)/useful life =40000/4 $10,000
Year 0 1 2 3 4
Initial Investment -$40,000
Operating cost -$29,000 -$29,000 -$29,000 -$29,000
Saving in cost $50,000 $50,000 $50,000 $50,000
Saving in Tax due to Dep. (27% of $10000) $2,700 $2,700 $2,700 $2,700
Total CashFlow -$40,000 $23,700 $23,700 $23,700 $23,700
PV Factor @ 9%= (1/(1.09^n)) 1 0.9174 0.8417 0.7722 0.7084
NPV -$40,000 $21,743 $19,948 $18,301 $16,790 $36,781
IRR (using IRR Formula) 46.33%
1c
DDB Dep. @50%
Year 1 2 3 4
Opening $40,000 $20,000 $10,000 $5,000
Dep. @ 50% $20,000 $10,000 $5,000 $5,000
Closing $20,000 $10,000 $5,000 $0
Year 0 1 2 3 4
Initial Investment -$40,000
Operating cost -$29,000 -$29,000 -$29,000 -$29,000
Saving in cost $50,000 $50,000 $50,000 $50,000
Saving in Tax due to Dep. (27% of Dep.) $5,400 $2,700 $1,350 $1,350
Total CashFlow -$40,000 $26,400 $23,700 $22,350 $22,350
PV Factor @ 9%= (1/(1.09^n)) 1 0.9174 0.8417 0.7722 0.7084
NPV -$40,000 $24,220 $19,948 $17,258 $15,833 $37,260
IRR (using IRR Formula) 48.40%

(Kindly Rate Positively, if found helpful. Thank You :))